OpenStudy (anonymous):

How do I rewrite 2x+3y=1,200 in Slope intercept form?

OpenStudy (jdoe0001):

if you were to solve say " 2x+3y=1,200 " for "y" what would that give you?

TheSmartOne (thesmartone):

Slope intercept form \(\Huge\color{blue}{y=mx+b}\)

OpenStudy (anonymous):

The equation for slope - intercept is y=mx+b. So y=2/3x+400 I believe.

OpenStudy (anonymous):

I don't really know what you mean jdoe0001

TheSmartOne (thesmartone):

@fluteprincess352 Take the 2x on the other side and dived by 3 so that way y is all alone...

TheSmartOne (thesmartone):

@typicalmatato You have to show and guide the person to the answer. Please take some time and read the Code of Conduct...

TheSmartOne (thesmartone):

@fluteprincess352 What have you gotten yet?

OpenStudy (anonymous):

sorry guys stepped out for a sec had to do something

TheSmartOne (thesmartone):

Sure no problem..

OpenStudy (anonymous):

so what do I do first?

TheSmartOne (thesmartone):

subtract 2x

TheSmartOne (thesmartone):

So what do you get now?

OpenStudy (anonymous):

do I subtract it from 3y or 1,200?

TheSmartOne (thesmartone):

2x+3y=1,200 -2x -2x ------------- 3y=-2x+1200

OpenStudy (anonymous):

I see thank you so much!

TheSmartOne (thesmartone):

Now all you have to do is divide by 3 on both sides

OpenStudy (anonymous):

oh so once I divide both sides by 3 I get 3y=-2x+400

OpenStudy (anonymous):

??

TheSmartOne (thesmartone):

No

TheSmartOne (thesmartone):

3y=-2x+1200 -- --- ---- 3 3 3 =y=-2/3x+400

TheSmartOne (thesmartone):

You have to divide everything by 3.

OpenStudy (anonymous):

oh ok I get it now thank you! :)

TheSmartOne (thesmartone):

And now your equation is in the format of y=mx+b and remember m is the slope and b is the y-intercept.

OpenStudy (anonymous):

yes I will remember this thank you!

TheSmartOne (thesmartone):

No problem and \(\Huge\color{purple}{Have~A~Great~Day!}\)

OpenStudy (anonymous):

you to! :)